The weights of a sample of n = 8 college men will be used to create a 95% confidence interval for the mean weight of all college men. What is the correct t* multiplier involved in calculating the interval?

A.

2.365

B.

2.306

C.

1.895

D.

1.96

E.

2.00
Answer #1

solution:

Degrees of freedom = df = n - 1 = 8 - 1 = 7

At 95% confidence level the t is ,

= 1 - 95% = 1 - 0.95 = 0.05

  / 2= 0.05 / 2 = 0.025

t /2,df = t0.025,7 =2.365 ( using student t table)

correct option a